OpenStudy (anonymous):

Solve the equation 43= 7w-15-5w w=???

OpenStudy (anonymous):

43= 7w-15-5w 43=(7-5)w -15 43=2w-15 43+15=2w-15+15 58=2w 58/2=w w=.....

OpenStudy (anonymous):

is it 58 div 2

OpenStudy (anonymous):

58/2=29

OpenStudy (anonymous):

Thanks